# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 4.3   Quark models
Table 4.1: Comparison of calculated masses of spin-$\nicefrac{{1}}{{2}}$ nucleon and spin-$\nicefrac{{3}}{{2}}$ $\Delta$ resonances with RPP values. The models are characterized by their interaction: One-gluon exchange (OGE) [548], Goldstone-boson exchange (GBE) [666], instanton-induced interactions (I.I.I.) without [549] or with additional Goldstone-boson exchange [671], the hypercentral constituent-quark model (HQM) [523], diquark (DQ) models [662, 521], and the algebraic model (AM) of Ref. [565]. n.c. stands for not calculated.
@@
[3] [548] [666] [549] [671] [523] [662] [521] [565]
  $N(1440)$ 1440$\pm$30 1540 1462 1518 1440 1463 1543 1511 1440
$N(1535)$ 1530$\pm$15 1460 1522 1435 1484 1524 1538 1537 1566
$\Delta(1600)$ 1570$\pm$70 1795 1725 1866 1589 1727 1714 1689 1646
$\Delta(1700)$ 1710$\pm$20 1620 1652 1628 1591 1573 1673 1625 1649
$N(1710)$ 1710$\pm$30 1770 1650 1729 1709 1752 1640 1766 1710
$N(1650)$ 1615$\pm$15 1535 1776 1660 1672 1692 1675 1625 1680
$N(1880)$ 1880$\pm$50 1880 n.c. 1950 1893 1828 n.c. 1890 1972
$N(1895)$ 1895$\pm$15 1945 n.c. 1901 1851 1861 n.c. 1888 1972
$\Delta(1920)$ 1920$\pm$50 1915 n.c. 1906 1894 1921 1930 2042 1909
$\Delta(1940)$ 1950$\pm$50 2080 n.c. 2100 1871 1900 2003 1910 1945
@@ 1 $N(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$ $1503(4)$ 0 $\nicefrac{{1}}{{4}}$ $1572$
$N(\mathbf{70},1^{-},\nicefrac{{3}}{{2}})$ $1690(24)$ 0 00 $1662$
$\Delta(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$ $1656(14)$ 0 00 $1662$
@@ 3 $N(\mathbf{56},1^{-},\nicefrac{{1}}{{2}})$ $1868(23)$ 1 0 $1858$
The relativistic Bonn model with instanton-induced interactions [549] solves the problem with the Roper resonance by lowering the baryon masses with a large component of scalar (good) diquarks; however, the $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ does not have scalar diquarks and still lies $296(70)$ MeV above its experimental value. Combining the short-range instanton interaction with long-range GBE results in a spectrum that shows reasonable agreement with the data [671], however at the expense of additional parameters. Finally, the approaches [523, 662, 521, 565] do not employ specific quark-quark interactions but use operators describing spin-spin and flavor correlations; here the differences between calculations and data are always below 100 MeV but no insight is gained into the internal dynamics of resonances.
@@
to fit the light baryon spectrum (c.o.g. masses) up to the second shell. The formula makes minimal assumptions and has five parameters: the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ mass, the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}-N(939)$ hyperfine splitting $\delta$, the level spacing $\omega$ whose coefficient is the shell number N, a confinement term with parameter $A$, and a parameter $D$ which is only active in the third shell. The values of $x_{\text{conf}}$ are extracted from Fig. 3.5 d. The effect of instanton-induced interaction is taken into account by a mass shift proportional to the fraction of ‘good’ diquarks in the wave function, which are the components highlighted in red in Fig. 3.1. This entails $I_{\text{sym}}=\frac{1}{2}$ for $\psi_{1}$, $I_{\text{sym}}=\frac{1}{4}$ for $\psi_{3}$ and zero for all other $\psi_{i}$, so this term only contributes to $N^{\ast}$ resonances with spin $S=\frac{1}{2}$ and the fully antisymmetric quark pair in $S$-wave. (Hence it is not applicable for the $N(\mathbf{56},1^{-},\nicefrac{{1}}{{2}})$ doublet in the third shell.) The square root is chosen to yield a better fit. Numerically, $\omega=1.3$ GeV2 and $A=1.6$ GeV2 are used, and the term $D=0.8$ GeV2 is added for the third shell. It is curious that the agreement with the data is even better than in the calculations discussed above; in particular, no flavor-dependent interaction term is required.
@@
So what have we learned about the baryon spectrum? Many quark models describe the spectrum reasonably well and several of them also the $Q^{2}$ dependence of various form factors and electroproduction amplitudes. In fact, different models describe the spectrum similarly well: one-gluon exchange, Goldstone-boson exchange, instanton-induced interactions, the hypercentral model, all capture the essential aspects of the light and strange baryon spectrum. This suggests that the spectrum is not overly sensitive to the specific form of the interactions between the quarks, except that flavor-dependent interactions appear necessary for the correct level ordering of baryons with opposite parity. Because adding more interaction terms like in Eq. (4.20) introduces new parameters, some information may also be absorbed in the model parameters when fitting the spectrum: The Capstick-Isgur model has 13 parameters [548] and the Bonn model 7–10 [549, 679], with similar numbers for the Goldstone-boson exchange and hypercentral models depending on the setup [670, 530]. Common challenges across the different models include the $\Lambda(1405)$, which indicates the absence of meson-baryon effects, and the missing resonances problem – while new experiments and improved partial-wave analyses have added new states to the RPP, the spectrum as of today is still much sparser than quark models predict.
@@
Given that the spectrum alone does not seem to suffice for discriminating between models, a more stringent test of the underlying dynamics are elastic and transition form factors. Here, relativity turns out to be crucial to get the correct magnitude and $Q^{2}$ dependence of the form factors. But this is also not sufficient, as the spacelike ($Q^{2}>0$) and timelike ($Q^{2}<0$) properties are intertwined [691, 692, 18, 693]. For instance, electromagnetic form factors exhibit vector-meson poles in the timelike region which significantly impact the spacelike properties, as is evident in dispersive calculations of form factors [691, 693, 694]. These effects must be dynamically generated from the quark level, which many quark models discussed above would not achieve. Is it then satisfactory if the model describes spacelike experimental data well, even if it does not produce timelike vector-meson poles whose effects are felt throughout the spacelike region? In other words, should one prioritize a pointwise description of spacelike data or genuine insights into the underlying dynamics? In this context one should highlight the covariant spectator model, which successfully connects the spacelike and timelike regions and has been explored for a range of elastic and transition form factors [695, 696, 697, 698, 699, 700], see [18] for a review.
@@
In general, the relativistic nature of light baryons tells us that neither non-relativistic nor relativistic quantum mechanics is sufficient for a complete description; instead, one must employ relativistic QFT. Unfortunately, translating the Schrödinger equation to relativistic QFT is not straightforward. Relativity forces us to abandon the concept of a wave function with a probability interpretation. It also calls the spin-flavor $SU(6)$ classification into question, because only the mass $M$ and spin $J$ are Casimirs of the Poincaré group, while $L$ and $S$ can mix in different frames and are no longer good quantum numbers for labeling states. Particle creation and annihilation makes the Hilbert space infinite-dimensional, and renormalization is highly non-trivial in a Hamiltonian description. In QFT, the spectrum can be accessed by extracting hadron poles from QCD’s correlation functions, which requires an understanding of how those correlation functions are generated to begin with. One way to deal with this is lattice QCD discussed in Sec. 4.6; however, its language is quite different from quark models which can make a meaningful dialogue difficult. Another option are the functional methods discussed in Sec. 4.5, which also address all of the above points but share sufficient similarities to provide a possible avenue for connecting quark models with QCD.
@@ 4.4   Light-front holography
A different approach to calculate the baryon spectrum is light-front holographic QCD [701, 702]. Here the idea is to convert the QCD Lagrangian into a light-front Hamiltonian $H_{\text{LF}}$ such that the spectrum and light-front wave functions are obtained from the eigenvalue equation $H_{\text{LF}}\,\psi=M^{2}\psi$ [703]. Because the diagonalization of this Hamiltonian is a difficult task, a practical simplification is to employ a semiclassical approximation and reduce the equation to a Schrödinger equation where the dynamics is encoded in an effective potential. It turns out that this equation has the same structure as the eigenvalue equations in AdS (Anti-de-Sitter) space, which is a spacetime with a constant negative curvature. This is the basis of light-front holographic models.
@@
The underlying idea is the holographic principle [704, 705, 706, 707], which is based on black hole thermodynamics and postulates that a gravitational system in a given volume is equivalent to a non-gravitational system on the boundary of that volume. A concrete realization of this principle is the AdS/CFT correspondence [708] between gravity on a $(d+1)$-dimensional AdS spacetime and a conformal field theory (CFT) living on the $d$-dimensional boundary of that spacetime. In particular, the correspondence implies that a strong coupling in the quantum gauge theory corresponds to a weak coupling in the classical gravitational theory (its holographic dual), which enables calculations of observables in nonperturbative regimes that are otherwise not accessible.
@@
Concerning QCD, however, so far no holographic dual is known. Among other reasons, QCD is not a conformal theory: Although the classical QCD Lagrangian for massless quarks is scale invariant, the regularization of the quantum field theory breaks this scale invariance. At low energies, QCD has mass gaps, confinement and dynamical chiral symmetry breaking, which are all hallmarks of a non-conformal theory. One strategy is then to employ a bottom-up approach named AdS/QCD, where one searches for a higher-dimensional classical gravitational theory that encodes basic aspects of QCD. Starting from a five-dimensional AdS background, one can break conformal symmetry e.g. by introducing an infrared cutoff (‘hard wall’ [709]) to mimic the confinement of hadronic modes in a box, or by a dilaton background (‘soft wall’ [710]) to model the running coupling or chiral symmetry breaking. Although not derived from first principles, this is a simple and flexible approach that captures qualitative features of confinement and the hadron spectrum. For example, the soft-wall model leads to the linear Regge trajectories observed in the hadron mass spectrum.
@@
In light-front holography, the dilaton profile in the AdS wave equation is determined by the mechanism proposed by de Alfaro, Fubini and Furlan [711], which also settles the form of the potential in the light-front Schrödinger equation. The resulting equation can be solved analytically to return the hadron masses and their light-front wave functions. For mesons, this leads to the mass formula
@@
where $J$ is the total angular momentum of the state, $L$ its orbital angular momentum and $n$ the radial quantum number. $\lambda$ is the coefficient of the dilaton field and plays the role of the dynamically generated mass scale in QCD, which can be traded for the $\rho$-meson mass $M_{\rho}$ given that the $\rho$ meson corresponds to $J=1$, $L=0$ and $n=0$. Note that Eq. (4.22) also implies linear Regge trajectories as well as a massless pion ($J=L=n=0$). Reviews of meson phenomenology using AdS/QCD can be found, e.g., in [712, 702, 713].
@@
where we traded the parameter $\lambda$ for the nucleon mass $M_{N}$, and $\nu$ is related to the minimal light-front orbital angular momentum of the baryon. Although this equation does not know about baryon multiplets, one can combine it with the nonrelativistic spin-flavor $SU(6)$ classification discussed in Sec. 3.2.1. Restricting ourselves to flavor octets and decuplets, the $\mathbf{56}$-plet contains nucleons with three-quark spin $S=\nicefrac{{1}}{{2}}$ and $\Delta$ baryons with $S=\nicefrac{{3}}{{2}}$, while the $\mathbf{70}$-plet consists of nucleons with $S=\nicefrac{{1}}{{2}}$ and $S=\nicefrac{{3}}{{2}}$ and $\Delta$ baryons with $S=\nicefrac{{1}}{{2}}$ (cf. Fig. 3.1). In Ref. [702] it is further assumed that the $\mathbf{56}$-plet only admits baryons with positive parity and even values of $L$, while the $\mathbf{70}$-plet only admits baryons with negative parity and odd values of $L$. This leads to a different scheme compared to Fig. 3.2, whose $\mathbf{70}$-plet entries with $L^{P}=0^{+}$ and $2^{+}$ would no longer appear here, so one also needs to reassign the physical states to different $SU(6)$ multiplets. The shell number N is also no longer available since the radial quantum number $n$ has taken its role. By comparing Regge trajectories, one is then led to the phenomenological identification $\nu=L+S/2-P/4$, where $P$ is the parity of the baryon, $S$ its intrinsic spin and $L$ its orbital angular momentum. Written out explicitly, this entails
$\begin{array}[]{ll}N(\mathbf{56},L^{+},\nicefrac{{1}}{{2}})&\leftrightarrow\quad\nu=L\,,\\ \Delta(\mathbf{56},L^{+},\nicefrac{{3}}{{2}})&\leftrightarrow\quad\nu=L+\frac{1}{2}\,,\end{array}\qquad\begin{array}[]{ll}\left\{\displaystyle N(\mathbf{70},L^{-},\nicefrac{{1}}{{2}})\atop\displaystyle\Delta(\mathbf{70},L^{-},\nicefrac{{1}}{{2}})\right\}&\leftrightarrow\quad\nu=L+\frac{1}{2}\,,\\ \;\;\;N(\mathbf{70},L^{-},\nicefrac{{3}}{{2}})&\leftrightarrow\quad\nu=L+1\,,\end{array}$ (4.28)
@@
Figure 4.10: Light baryon spectrum in the light-front holographic model discussed in Ref. [702]. The left panel shows the positive-parity spectrum, the center panel the negative-parity spectrum and the right panel the corresponding Regge trajectories.
@@
with $\kappa=0$, $1/2$ or $1$ for the three respective options in Eq. (4.29).
@@
In general, light-front holography is a powerful tool to provide analytic insight into QCD phenomenology. As discussed above, it naturally produces Regge trajectories and gives overall good agreement with experimental data. Remarkably, the masses of baryons, the Regge slope and the hyperfine splitting are all described by one parameter, pointing at an intimate connection between these quantities. Regarding hadron phenomenology, apart from the baryon spectrum light-front holography has been applied to calculate distribution amplitudes, form factors, electromagnetic transition amplitudes and a range of other quantities, see e.g. [709, 714, 715, 716, 717, 718, 719, 720, 721, 712, 722, 702, 723, 2] and references therein.

### v4

@@ 4.3  Quark models
Table 4.1: Comparison of calculated masses of spin-$\nicefrac{{1}}{{2}}$ nucleon and spin-$\nicefrac{{3}}{{2}}$ $\Delta$ resonances with RPP values. The models are characterized by their interaction: One-gluon exchange (OGE) Capstick:1986ter , Goldstone-boson exchange (GBE) Glozman:1997ag , instanton-induced interactions (I.I.I.) without Loring:2001kx or with additional Goldstone-boson exchange Ronniger:2012xp , the hypercentral constituent-quark model (HQM) Giannini:2001kb , diquark (DQ) models Santopinto:2004hw , Santopinto:2014opa , and the algebraic model (AM) of Ref. Bijker:1994yr . n.c. stands for not calculated.
@@
ParticleDataGroup:2024cfk Capstick:1986ter Glozman:1997ag Loring:2001kx Ronniger:2012xp Giannini:2001kb Santopinto:2004hw Santopinto:2014opa Bijker:1994yr
[…]
@@ 1 $N(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$ $1503(4)$ 0 $\nicefrac{{1}}{{4}}$ $1572$
$N(\mathbf{70},1^{-},\nicefrac{{3}}{{2}})$ $1690(24)$ 0 0 $1662$
$\Delta(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$ $1656(14)$ 0 0 $1662$
@@ 3 $N(\mathbf{56},1^{-},\nicefrac{{1}}{{2}})$ $1868(23)$ 1 0 $1858$
The relativistic Bonn model with instanton-induced interactions Loring:2001kx solves the problem with the Roper resonance by lowering the baryon masses with a large component of scalar (good) diquarks; however, the $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ does not have scalar diquarks and still lies $296(70)$ MeV above its experimental value. Combining the short-range instanton interaction with long-range GBE results in a spectrum that shows reasonable agreement with the data Ronniger:2012xp , however at the expense of additional parameters. Finally, the approaches Giannini:2001kb , Santopinto:2004hw , Santopinto:2014opa , Bijker:1994yr do not employ specific quark-quark interactions but use operators describing spin-spin and flavor correlations; here the differences between calculations and data are always below 100 MeV but no insight is gained into the internal dynamics of resonances.
@@
to fit the light baryon spectrum (c.o.g. masses) up to the third shell. The formula makes minimal assumptions and has five parameters: the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ mass, the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}-N(939)$ hyperfine splitting $\delta$, the level spacing $\omega$ whose coefficient is the shell number N, a confinement term with parameter $A$, and a parameter $D$ which is only active in the third shell. The values of $x_{\text{conf}}$ are extracted from Fig. 3.5 d. The effect of instanton-induced interaction is taken into account by a mass shift proportional to the fraction of ‘good’ diquarks in the wave function, which are the components highlighted in red in Fig. 3.1. This entails $I_{\text{sym}}=\frac{1}{2}$ for $\psi_{1}$, $I_{\text{sym}}=\frac{1}{4}$ for $\psi_{3}$ and zero for all other $\psi_{i}$, so this term only contributes to $N^{\ast}$ resonances with spin $S=\frac{1}{2}$ and the fully antisymmetric quark pair in $S$-wave. (Hence it is not applicable for the $N(\mathbf{56},1^{-},\nicefrac{{1}}{{2}})$ doublet in the third shell.) The square root is chosen to yield a better fit. Numerically, $\omega=1.3$ GeV2 and $A=1.6$ GeV2 are used, and the term $D=0.8$ GeV2 is added for the third shell. It is curious that the agreement with the data is even better than in the calculations discussed above; in particular, no flavor-dependent interaction term is required.
@@
So what have we learned about the baryon spectrum? Many quark models describe the spectrum reasonably well and several of them also the $Q^{2}$ dependence of various form factors and electroproduction amplitudes. In fact, different models describe the spectrum similarly well: one-gluon exchange, Goldstone-boson exchange, instanton-induced interactions, the hypercentral model, all capture the essential aspects of the light and strange baryon spectrum. This suggests that the spectrum is not overly sensitive to the specific form of the interactions between the quarks, except that flavor-dependent interactions appear necessary for the correct level ordering of baryons with opposite parity. Because adding more interaction terms like in Eq. (4.20) introduces new parameters, some information may also be absorbed in the model parameters when fitting the spectrum: The Capstick-Isgur model has 13 parameters Capstick:1986ter and the Bonn model 7–10 Loring:2001kx , Ronniger:2011td , with similar numbers for the Goldstone-boson exchange and hypercentral models depending on the setup Plessas:2015mpa , Giannini:2015zia . Common challenges across the different models include the $\Lambda(1405)$, which indicates the absence of meson-baryon effects, and the missing resonances problem – while new experiments and improved partial-wave analyses have added new states to the RPP, the spectrum as of today is still much sparser than quark models predict.
@@
Given that the spectrum alone does not seem to suffice for discriminating between models, a more stringent test of the underlying dynamics are elastic and transition form factors. Here, relativity turns out to be crucial to get the correct magnitude and $Q^{2}$ dependence of the form factors. But this is also not sufficient, as the spacelike ($Q^{2}>0$) and timelike ($Q^{2}<0$) properties are intertwined Denig:2012by , Eichmann:2016yit , Ramalho:2023hqd , Lin:2021umz . For instance, electromagnetic form factors exhibit vector-meson poles in the timelike region which significantly impact the spacelike properties, as is evident in dispersive calculations of form factors Denig:2012by , Lin:2021umz , Alvarado:2023loi . These effects must be dynamically generated from the quark level, which many quark models discussed above would not achieve. Is it then satisfactory if the model describes spacelike experimental data well, even if it does not produce timelike vector-meson poles whose effects are felt throughout the spacelike region? In other words, should one prioritize a pointwise description of spacelike data or genuine insights into the underlying dynamics? In this context one should highlight the covariant spectator model, which successfully connects the spacelike and timelike regions and has been explored for a range of elastic and transition form factors Gross:2006fg , Ramalho:2008dp , Ramalho:2012ng , Ramalho:2015qna , Ramalho:2016zgc , Ramalho:2020nwk , see Ramalho:2023hqd for a review.
@@
In general, the relativistic nature of light baryons tells us that neither non-relativistic nor relativistic quantum mechanics is sufficient for a complete description; instead, one must employ relativistic QFT. Unfortunately, translating the Schrödinger equation to relativistic QFT is not straightforward. Special relativity forces us to abandon the concept of a wave function with a probability interpretation. It also calls the spin-flavor $SU(6)$ classification into question, because only the mass $M$ and spin $J$ are Casimirs of the Poincaré group, while $L$ and $S$ can mix in different frames and are no longer good quantum numbers for labeling states. Particle creation and annihilation makes the Hilbert space infinite-dimensional, and renormalization is highly non-trivial in a Hamiltonian description. In QFT, the spectrum can be accessed by extracting hadron poles from QCD’s correlation functions, which requires an understanding of how those correlation functions are generated to begin with. One way to deal with this is lattice QCD discussed in Sec. 4.6; however, its language is quite different from quark models which can make a meaningful dialogue difficult. Another option are the functional methods discussed in Sec. 4.5, which also address all of the above points but share sufficient similarities to provide a possible avenue for connecting quark models with QCD.
@@ 4.4  Light-front holography
A different approach to calculate the baryon spectrum is light-front holographic QCD Brodsky:2008pg , Brodsky:2014yha . Here the idea is to convert the QCD Lagrangian into a light-front Hamiltonian $H_{\text{LF}}$ such that the spectrum and light-front wave functions are obtained from the eigenvalue equation $H_{\text{LF}}\,\psi=M^{2}\psi$ Brodsky:1997de . Because the diagonalization of this Hamiltonian is a difficult task, a practical simplification is to employ a semiclassical approximation and reduce the equation to a Schrödinger equation where the dynamics is encoded in an effective potential. It turns out that this equation has the same structure as the eigenvalue equations in AdS (Anti-de-Sitter) space, which is a spacetime with a constant negative curvature. This is the basis of light-front holographic models.
@@
The underlying idea is the holographic principle Bekenstein:1973ur , Hawking:1975vcx , tHooft:1993dmi , Susskind:1994vu , which is based on black hole thermodynamics and postulates that a gravitational system in a given volume is equivalent to a non-gravitational system on the boundary of that volume. A concrete realization of this principle is the AdS/CFT correspondence Maldacena:1997re between gravity on a $(d+1)$-dimensional AdS spacetime and a conformal field theory (CFT) living on the $d$-dimensional boundary of that spacetime. In particular, the correspondence implies that a strong coupling in the quantum gauge theory corresponds to a weak coupling in the classical gravitational theory (its holographic dual), which enables calculations of observables in nonperturbative regimes that are otherwise not accessible.
@@
Concerning QCD, however, so far no holographic dual is known. Among other reasons, QCD is not a conformal theory: Although the classical QCD Lagrangian for massless quarks is scale invariant, the regularization of the quantum field theory breaks this scale invariance. At low energies, QCD has mass gaps, confinement and dynamical chiral symmetry breaking, which are all hallmarks of a non-conformal theory. One strategy is then to employ a bottom-up approach named AdS/QCD, where one searches for a higher-dimensional classical gravitational theory that encodes basic aspects of QCD. Starting from a five-dimensional AdS background, one can break conformal symmetry e.g. by introducing an infrared cutoff (‘hard wall’ Polchinski:2001tt ) to mimic the confinement of hadronic modes in a box, or by a dilaton background (‘soft wall’ Karch:2006pv ) to model the running coupling or chiral symmetry breaking. Although not derived from first principles, this is a simple and flexible approach that captures qualitative features of confinement and the hadron spectrum. For example, the soft-wall model leads to the linear Regge trajectories observed in the hadron mass spectrum.
@@
In light-front holography, the dilaton profile in the AdS wave equation is determined by the mechanism proposed by de Alfaro, Fubini and Furlan deAlfaro:1976vlx , which also settles the form of the potential in the light-front Schrödinger equation. The resulting equation can be solved analytically to return the hadron masses and their light-front wave functions. For mesons, this leads to the mass formula
@@
where $J$ is the total angular momentum of the state, $L$ its orbital angular momentum and $n$ the radial quantum number. $\lambda$ is the coefficient of the dilaton field and plays the role of the dynamically generated mass scale in QCD, which can be traded for the $\rho$-meson mass $M_{\rho}$ given that the $\rho$ meson corresponds to $J=1$, $L=0$ and $n=0$. Note that Eq. (4.22) also implies linear Regge trajectories as well as a massless pion ($J=L=n=0$). Reviews of meson phenomenology using AdS/QCD can be found, e.g., in deTeramond:2011aml , Brodsky:2014yha , Kim:2011ey .
@@
where we traded the parameter $\lambda$ for the nucleon mass $M_{N}$, and $\nu$ is related to the minimal light-front orbital angular momentum of the baryon. Although this equation does not know about baryon multiplets, one can combine it with the nonrelativistic spin-flavor $SU(6)$ classification discussed in Sec. 3.2.1. Restricting ourselves to flavor octets and decuplets, the $\mathbf{56}$-plet contains nucleons with three-quark spin $S=\nicefrac{{1}}{{2}}$ and $\Delta$ baryons with $S=\nicefrac{{3}}{{2}}$, while the $\mathbf{70}$-plet consists of nucleons with $S=\nicefrac{{1}}{{2}}$ and $S=\nicefrac{{3}}{{2}}$ and $\Delta$ baryons with $S=\nicefrac{{1}}{{2}}$ (cf. Fig. 3.1). In Ref. Brodsky:2014yha it is further assumed that the $\mathbf{56}$-plet only admits baryons with positive parity and even values of $L$, while the $\mathbf{70}$-plet only admits baryons with negative parity and odd values of $L$. This leads to a different scheme compared to Fig. 3.2, whose $\mathbf{70}$-plet entries with $L^{P}=0^{+}$ and $2^{+}$ would no longer appear here, so one also needs to reassign the physical states to different $SU(6)$ multiplets. The shell number N is also no longer available since the radial quantum number $n$ has taken its role. By comparing Regge trajectories, one is then led to the phenomenological identification $\nu=L+S/2-P/4$, where $P$ is the parity of the baryon, $S$ its intrinsic spin and $L$ its orbital angular momentum. Written out explicitly, this entails
[…]
@@
Figure 4.10: Light baryon spectrum in the light-front holographic model discussed in Ref. Brodsky:2014yha . The left panel shows the positive-parity spectrum, the center panel the negative-parity spectrum and the right panel the corresponding Regge trajectories.
@@
with $\kappa=0$, $1/2$ or $1$ for the three respective options in Eq. (4.29). The connection between AdS/QCD and the non-relativistic three-oscillator model is, however, a speculation and not covered by the holographic approach to the spectrum Guy:2025pc .
@@
In general, light-front holography is a powerful tool to provide analytic insight into QCD phenomenology. As discussed above, it naturally produces Regge trajectories and gives overall good agreement with experimental data. Remarkably, the masses of baryons, the Regge slope and the hyperfine splitting are all described by one parameter, pointing at an intimate connection between these quantities. Regarding hadron phenomenology, apart from the baryon spectrum light-front holography has been applied to calculate distribution amplitudes, form factors, electromagnetic transition amplitudes and a range of other quantities, see e.g. Polchinski:2001tt , Katz:2005ir , Brodsky:2006uqa , Boschi-Filho:2005xct , Brodsky:2007hb , Forkel:2007cm , Brodsky:2008pf , Colangelo:2008us , Branz:2010ub , deTeramond:2011aml , Li:2013oda , Brodsky:2014yha , Leutgeb:2019gbz , Aliberti:2025beg and references therein.